OpenStudy (anonymous):

Geometry The sum of the measures of angles x and y is 180. the measure of angles x is 24 greater than the measure of angle y. a. define the variables, and write equations for this situtation b.find the measure of each angle

OpenStudy (anonymous):

x and y are angles. x+y=180, x-y=24 (in degrees). When you have a word problem, start by trying to write it in numbers and symbols so you can be sure to understand the problem as a math problem, and (because problems given in textbooks and lectures are usually intended to be level-appropriate) it's usually quite simple to solve the problem once you've found equations.

OpenStudy (goformit100):

@Numb3r1 it is correct